If the temperature of a sample of methane gas rises by 13 ∘C when 8.8 kJ of heat energy is added to the sample, what is the mass of the sample?
1
Expert's answer
2018-01-22T05:55:22-0500
Cp (CH4) = 2483 J/(kg*K)
T = 13 oC
E = 8800 J
m = E/(Cp*T)
m = 8800/(2483*13) = 0.273(g(CH4)